# Variable Substitution and Bilinear Programming for Aligning Partially Overlapping Point Sets

### v1

@@ 5 Optimization
In this section, we begin by applying the bilinear relaxation formula, introduced in Section 3, to derive a lower bound function for our objective. This lower bound function’s optimization can then be divided into separate optimizations over $\boldsymbol{\theta}$ and $\mathbf{p}$ (discussed in Section 5.1). Subsequently, we demonstrate that it suffices to branch solely over $\boldsymbol{\theta}$ (explained in Section 5.2) and introduce a more effective method for computing the upper bound (outlined in Section 5.4). Lastly, we introduce the BnB algorithm (presented in Section 5.5), which offers global optimization for the objective function.
@@ 5.1 Lower bound function derivation
We note that $E$ contains bilinear terms $\text{trace}(\boldsymbol{\Gamma}\boldsymbol{\Theta})$ and $\boldsymbol{\theta}^{\top}\boldsymbol{\eta}$, which interconnect $\boldsymbol{\Gamma}$ and $\boldsymbol{\Theta}$ (resp. $\boldsymbol{\theta}$ and $\boldsymbol{\eta}$), making optimization challenging. To tackle this challenge, we use the bilinear relaxation formula introduced in Section 3 to relax these terms into linear ones. The first relaxation is:
@@
where in accordance with equations (1) and (2), the coefficients can be calculated as
@@
Here we use underline (resp. overline) to denote the lower (resp. upper) bounds of matrices $\boldsymbol{\Theta}$ and $\boldsymbol{\Gamma}$. It is also worth noting that the range $[\underline{\boldsymbol{\Theta}}$, $\overline{\boldsymbol{\Theta}}]$ of $\boldsymbol{\Theta}$ can be derived from the range $[\underline{\boldsymbol{\theta}},\overline{\boldsymbol{\theta}}]$ of $\boldsymbol{\theta}$ based on relation $\boldsymbol{\Theta}=\boldsymbol{\theta}\boldsymbol{\theta}^{\top}$, as explained in Section 5.3.
@@
where in accordance with equations (1) and (2), the coefficients can be calculated as
@@
By substituting back $\boldsymbol{\theta}\boldsymbol{\theta}^{\top}\Leftarrow\boldsymbol{\Theta}$, $\text{mat}(K\mathbf{B}_{2}\mathbf{p})\Leftarrow\boldsymbol{\Gamma}$ and $-2\mathbf{A}\mathbf{p}\Leftarrow\boldsymbol{\eta}$, we get the final form of the lower bound function:
@@
The optimization over $\mathbf{p}$ is
@@
The optimization over $\boldsymbol{\theta}$ is:
@@ 5.2 Branching strategy
According to the result stated in Section 3, for the bilinear term $\boldsymbol{\theta}^{\top}\boldsymbol{\eta}$, for its linear lower bound function to converge to it, it suffices to branch over $\boldsymbol{\theta}$ and leave the range of $\boldsymbol{\eta}$ fixed. Here, utilizing $\boldsymbol{\eta}=-2\mathbf{A}\mathbf{p}$, the fixed range $[\underline{\boldsymbol{\eta}},\overline{\boldsymbol{\eta}}]$ of $\boldsymbol{\eta}$ can be computed as
$\underline{\eta}_{i}=\min_{\mathbf{p}\in\Omega}(-2\mathbf{A}\mathbf{p})_{i}$ (15)
$\overline{\eta}_{i}=\max_{\mathbf{p}\in\Omega}(-2\mathbf{A}\mathbf{p})_{i}$ (16)
Likewise, for the bilinear term $\text{trace}(\boldsymbol{\Gamma}\boldsymbol{\Theta})$, for its linear lower bound function to converge to it, we only need to branch over $\boldsymbol{\theta}$ (since the range of $\boldsymbol{\theta}$ determines the range of $\boldsymbol{\Theta}$ as explained in Sec. 5.3) and leave the range of $\boldsymbol{\Gamma}$ fixed. Here, utilizing $\boldsymbol{\Gamma}=\text{mat}(K\mathbf{B}_{2}\mathbf{p})$, the fixed range $[\underline{\boldsymbol{\Gamma}},\overline{\boldsymbol{\Gamma}}]$ of $\boldsymbol{\Gamma}$ can be computed as follows:
$\underline{\Gamma}_{ij}=\min_{\mathbf{p}\in\Omega}[\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})]_{ij}$ (17)
$\overline{\Gamma}_{ij}=\max_{\mathbf{p}\in\Omega}[\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})]_{ij}$ (18)
With thess choice of $\underline{\boldsymbol{\Gamma}}$ and $\overline{\boldsymbol{\Gamma}}$, it has been empirically demonstrated that
@@
Please refer to Section 5.6 for relevant experiments and Appendix for a weaker assertion. This result implies that problem (14a) (14b) is now a convex quadratic program. Consequently, utilizing solvers for convex quadratic programs, such as MATLAB’s quadprog function, guarantees global optimality. We will address the proof of (19) in our future research.
We also note that (15), (16), (17) and (18) are $n_{p}-$cardinlaity linear assignment problems, which can be efficiently solved by the aforementioned algorithms.
@@ 5.3 The range of $\Theta$ from the range of $\theta$
$[\underline{\Theta_{ij}},\overline{\Theta_{ij}}]=\begin{cases}[\underline{\theta_{i}},\overline{\theta_{i}}]\cdot[\underline{\theta_{j}},\overline{\theta_{j}}],&\text{if}\ i\neq j\\ [\underline{\theta_{i}},\overline{\theta_{i}}]^{2},&\text{if}\ i=j\end{cases}$
@@
$[\underline{\theta_{i}},\overline{\theta_{i}}]\cdot[\underline{\theta_{j}},\overline{\theta_{j}}]=[\min(\underline{\theta_{i}}\underline{\theta_{j}},\underline{\theta_{i}}\overline{\theta_{j}},\overline{\theta_{i}}\underline{\theta_{j}},\overline{\theta_{i}}\overline{\theta_{j}}),\max(\underline{\theta_{i}}\underline{\theta_{j}},\underline{\theta_{i}}\overline{\theta_{j}},\overline{\theta_{i}}\underline{\theta_{j}},\overline{\theta_{i}}\overline{\theta_{j}})]$
@@
$[\underline{\theta_{i}},\overline{\theta_{i}}]^{2}=\begin{cases}[\min(\underline{\theta_{i}}^{2},\overline{\theta_{i}}^{2}),\max(\underline{\theta_{i}}^{2},\overline{\theta_{i}}^{2})],&\text{if}\ 0\notin[\underline{\theta_{i}},\overline{\theta_{i}}]\\ [0,\max(\underline{\theta_{i}}^{2},\overline{\theta_{i}}^{2})],&\text{if}\ 0\in[\underline{\theta_{i}},\overline{\theta_{i}}]\end{cases}$
@@ 5.5 Branch-and-Bound
Based on the aforementioned preparations, we are now poised to apply the BnB algorithm to optimize $E$. To initiate the process, we utilize the initial range of $\boldsymbol{\theta}$ to form the initial hypercube $M$. Subsequently, during each iteration of the algorithm, we identify the hypercube that yields the lowest lower bound among all available hypercubes. This selected hypercube is then further subdivided to enhance the global lower bound of the problem. Simultaneously, the upper bound is refined by assessing $E(\mathbf{p})$ using solutions generated during the computation of the lower bounds. For a comprehensive overview, refer to Algorithm 1.
@@ 5.6 Positive semidefiniteness of $\mathbf{H}^{0}+\mathbf{C}$
To confirm the assertion that $\mathbf{H}^{0}+\mathbf{C}\succeq 0$, we utilize the separate outliers and inliers test detailed in Section 8.1.1, with an outlier to data ratio set at 0.3. We present histograms of the smallest eigenvalues of the matrix $\mathbf{H}^{0}+\mathbf{C}$ across 100 registration instances in Figure 3. The result consistently support the claim that $\mathbf{H}^{0}+\mathbf{C}\succeq 0$, underscoring the efficacy of the proposed algorithm.
@@
5.7 Convergence of the BnB algorithm
To assess the convergence of the proposed BnB algorithm, we conducted the separate outliers and inliers test detailed in Sec. 8.1.1, with an outlier to data ratio set at $0.3$. The experimental results are depicted in Fig. 4. Key observations include: 1) A narrower initial range of $\boldsymbol{\theta}$ results in a tighter duality gap, evidenced by improved upper and lower bounds. This suggests that leveraging additional information about the range of $\boldsymbol{\theta}$ could effectively reduce the duality gap in practical applications. 2) Easier problem instances, such as the fish test compared to the character test, exhibit a tighter duality gap, thus leads to faster convergence.
Properties 2) suggests that employing a fixed duality gap threshold as the stopping criterion for all types of problems may not be a good idea. It may happen that a threshold is set too small for challenging problems, leading to prolonged execution times. Conversely, it might be set too large for simpler problems, resulting in premature termination. Therefore, following [6], we adopt the maximum branching depth as our stopping criterion.
Figure 4: Upper (first row) and lower bounds (second row) generated in each iteration of RPM-BP. Our method is tested with the $n_{p}$ value chosen as the ground truth and with varying initial ranges of $\boldsymbol{\theta}$: $[\boldsymbol{\theta}_{gt}-\Delta,\boldsymbol{\theta}_{gt}+\Delta]$, where $\boldsymbol{\theta}_{gt}$ represents the ground truth $\boldsymbol{\theta}$ solution. Here, the margin $\Delta$ takes values of $0.5$, $1$, and $1.5$ respectively.
6 Application 1: 2D similarity/affine registration
Our algorithm can handle 2D similarity and affine transformations since both of them can be written as the form of $T(\mathbf{x}|\boldsymbol{\theta})=J(\mathbf{x})\boldsymbol{\theta}$. A 2D similarity transformation has the form
$T(\mathbf{x}|{\boldsymbol{\theta}})=\begin{bmatrix}\theta_{1}&-\theta_{2}\\ \theta_{2}&\theta_{1}\end{bmatrix}\begin{bmatrix}x^{1}\\ x^{2}\end{bmatrix}+\begin{bmatrix}\theta_{3}\\ \theta_{4}\end{bmatrix}=\begin{bmatrix}x^{1}&-x^{2}&1&0\\ x^{2}&x^{1}&0&1\end{bmatrix}\boldsymbol{\theta}$
where $[\theta_{3},\theta_{4}]^{\top}$ is translation and $\theta_{1}=s\cos(\phi)$ and $\theta_{2}=s\sin(\phi)$ with $s$ being scale and $\phi$ being rotation angle. It can be derived that $\mathbf{B}_{2}=\mathbf{B}([1,3,4],:)$ and $\mathbf{C}=\text{diag}([0,0,n_{p},n_{p}])$.
A 2D affine transformation has the form
$T(\mathbf{x}|{\boldsymbol{\theta}})=\begin{bmatrix}\theta_{1}&\theta_{2}\\ \theta_{3}&\theta_{4}\end{bmatrix}\begin{bmatrix}x^{1}\\ x^{2}\end{bmatrix}+\begin{bmatrix}\theta_{5}\\ \theta_{6}\end{bmatrix}=\begin{bmatrix}x^{1}&x^{2}&0&0&1&0\\ 0&0&x^{1}&x^{2}&0&1\end{bmatrix}\boldsymbol{\theta}$
It can be derived that $\mathbf{B}_{2}=\mathbf{B}([1,2,5,8,11],:)$ and $\mathbf{C}=\text{diag}([0,0,0,0,n_{p},n_{p}])$.
7 Application 2: 3D rigid registration
Theoretically, our method can handle 3D affine transformation since it takes the form $T(\mathbf{x}|\boldsymbol{\theta})=J(\mathbf{x})\boldsymbol{\theta}$. A 2D affine transformation has the form
$T(\mathbf{x}|{\boldsymbol{\theta}})=\begin{bmatrix}\theta_{1}&\theta_{2}&\theta_{3}\\ \theta_{4}&\theta_{5}&\theta_{6}\\ \theta_{7}&\theta_{8}&\theta_{9}\end{bmatrix}\begin{bmatrix}x^{1}\\ x^{2}\\ x^{3}\end{bmatrix}+\begin{bmatrix}\theta_{10}\\ \theta_{11}\\ \theta_{12}\end{bmatrix}$
=$\begin{bmatrix}x^{1}&x^{2}&x^{3}&0&0&0&0&0&0&1&0&0\\ 0&0&0&x^{1}&x^{2}&x^{3}&0&0&0&0&1&0\\ 0&0&0&0&0&0&x^{1}&x^{2}&x^{3}&0&0&1\end{bmatrix}\boldsymbol{\theta}$
It can be derived that $\mathbf{B}_{2}=\mathbf{B}([1,2,3,10,14,15,22,27,34],:)$ and $\mathbf{C}=\text{diag}([\underbrace{0,\ldots,0}_{9\text{ zeros}},n_{p},n_{p},n_{p}])$. Nevertheless, the 3D affine transformation has many parameters, causing our method to converge slowly. Therefore, for the 3D case, we instead focus on rigid transformation instead.
[…]
where $[\cdot]_{\times}$ denotes the skew-symmetric matrix representation:
where $r_{i}$ is the $i$-th element of $\mathbf{r}$.
3D rigid transformation is a special case of 3D affine transformation, whereas our method is capable of handling 3D affine transformation. Therefore, we aim to adapt our method from Section 5 to accommodate 3D rigid transformation as well. Here’s the procedure: We continue to use the algorithm outlined in Section 5 for registration, where the transformation is defined as a 3D affine transformation: $T(\mathbf{x}|\mathbf{R},\mathbf{t})=\mathbf{R}\mathbf{x}+\mathbf{t}$, with $\mathbf{R}$ representing the linear part of the affine transformation. The main differences from the algorithm in Section 5 are as follows: (1) Instead of branching over $\mathbf{R}$ and $\mathbf{t}$, we branch over $\mathbf{r}$ and $\mathbf{t}$. (2) In each iteration of the BnB algorithm, given the range of $\mathbf{r}$, we calculate the range of $\mathbf{R}$ from the range of $\mathbf{r}$ based on equation (21). (The method for efficiently computing the range of $\mathbf{R}$ is discussed in Section 7.1.)
In the above algorithm, (20) is employed to compute the upper bound, assuming that the transformation can be affine. While methods exist (see Sec. 5.3 of [6]) for computing the upper bound with a rigid transformation constraint, they often result in higher upper bounds, which slows down our method’s convergence. Hence, we opt to use (20) for computing the upper bound in this work.
7.1 The range of $\mathbf{R}$ from the range of $\mathbf{r}$
Given a range $[\underline{\mathbf{r}},\overline{\mathbf{r}}]$ for $\mathbf{r}$, the computation of the range of $\mathbf{R}$, as derived from (21), involves determining $R_{ij}$ values within the bounds specified by $\underline{\mathbf{r}}$ and $\overline{\mathbf{r}}$. This computation can be facilitated through optimization solvers such as the Matlab function $fmincon$. However, this process can be burdensome, particularly when executed iteratively as a subroutine of the Branch and Bound (BnB) algorithm. In order to alleviate this computational burden, we propose the utilization of precomputation techniques to enhance efficiency. Specifically, prior to initiating the BnB algorithm, we construct a regular grid, with a predefined width (set as $150$ in this study), over the initial cube $[\underline{\mathbf{r}}_{0},\overline{\mathbf{r}}_{0}]$. Subsequently, we compute the corresponding $R_{ij}$ values for all grid points and store them. During the execution of the BnB algorithm, when presented with a range $[\underline{\mathbf{r}},\overline{\mathbf{r}}]$, we first identify the grid points falling within this range. Subsequently, we determine the minimum and maximum precomputed $R_{ij}$ values for these points, effectively approximating $\min\{R_{ij}|\underline{\mathbf{r}}\leq\mathbf{r}\leq\overline{\mathbf{r}}\}$ and $\max\{R_{ij}|\underline{\mathbf{r}}\leq\mathbf{r}\leq\overline{\mathbf{r}}\}$.
8 Experiments
We implement the proposed algorithm, referred to as RPM-BP, using Matlab R2023b. To evaluate its performance, we conduct comparisons with other methods on a computer equipped with 6-core CPUs running at 3.2 GHz.
For the competing methods that solely produce point correspondences, we utilize the generated point correspondences to compute the optimal affine transformation between two sets of points. The matching error is then quantified as the root-mean-square distance between the transformed model inliers and their corresponding scene inliers.
To ensure efficient computation, we set the maximum branching depth of RPM-BP to 12.
8.1 2D registration
We conduct comparisons with RPM-HTB [6], RPM-PA [5] and RPM-CAV [4], all of which utilize global optimization techniques. These methods are adept at handling partial overlap and permit arbitrary similarity transformations between two point sets, rendering them suitable candidates for our comparative analysis.
8.1.1 2D synthetic data
Synthetic data offer a convenient means to assess specific facets of an algorithm’s performance. We conduct five distinct types of tests to gauge a method’s resilience to various disturbances: i) Deformation test: The prototype shape undergoes non-rigid deformation to generate the scene set. ii) Noise test: Positional noise is introduced to perturb the prototype shape, resulting in the scene point set. iii) Mixed outliers and inliers test: Random outliers are superimposed on the prototype shape to create both point sets. iv) Separate outliers and inliers test: Random outliers are added to different areas of the prototype shape, generating the two point sets. v) Occlusion+outlier test: Initially, the prototype shape is occluded to produce the two point sets. Subsequently, random outliers (with a fixed outlier-to-data ratio of 0.5) are incorporated into different regions of the point sets.
These tests are visually illustrated in Fig. 7. Additionally, to assess a method’s capability in handling arbitrary rotation and scaling, random rotations and scalings within the range of $[0.5,1.5]$ are applied when generating the two point sets. Fig. 7 provides examples of registrations performed by different methods.
The registration errors produced by various methods are depicted in the top two rows of Fig. 7. The result suggests that, in comparison to alternative methods, RPM-BP and RPM-HTB, with $n_{p}$ chosen as the ground truth, exhibits robustness against deformation, noise, and outliers when the outliers are distinct from the inliers. However, their robustness diminishes when outliers are intermixed with inliers. This limitation arises due to the fact that our method lacks $\epsilon-$global optimality, as discussed in Sec. 5.7, thereby affecting its robustness.
Regarding different transformation options, RPM-BP, when employing similarity transformation, demonstrates greater resilience to various disturbances than when utilizing affine transformation.
Furthermore, the choice of $n_{p}$ value significantly influences the performance of RPM-BP. Specifically, RPM-BP with an $n_{p}$ value close to the ground truth yields superior results. Moreover, RPM-BP displays less sensitivity to variations in the $n_{p}$ value compared to RPM-HTB, particularly evident in tests involving outliers and occlusion.
The average run times for various methods are depicted in the bottom row of Fig. 7. RPM-BP exhibits the best performance in scenarios involving deformation, positional noise, and mixed inlier-outlier tests. Comparing different transformation options, RPM-BP employing similarity transformation proves to be more efficient than when employing affine transformation. This efficiency is attributed to the lesser number of parameters involved in similarity transformation.
The outlier tests also provide insights into how various methods scale with problem size. Among them, RPM-BP, RPM-HTB, and RPM-CAV demonstrate the best scalability, followed by RPM-PA.
Refer to caption Refer to caption Refer to caption

### v2

@@ 5 Optimization
This section details the optimization strategy for our Branch-and-Bound (BnB) algorithm. We begin by deriving a lower bound for our objective function using bilinear relaxation (Sec. 5.1). We show that the minimization of this lower bound can be efficiently decoupled into two independent subproblems. Next, we present our branching strategy (Sec. 5.2), demonstrating that it’s sufficient to branch on only the low-dimensional transformation parameters. Finally, we discuss how to compute a more effective upper bound (Sec. 5.4) and outline the complete BnB algorithm (Sec. 5.5).
@@ 5.1 Lower bound function derivation
We note that $E$ contains bilinear terms $\text{trace}(\boldsymbol{\Gamma}\boldsymbol{\Theta})$ and $\boldsymbol{\theta}^{\top}\boldsymbol{\eta}$, which interconnect $\boldsymbol{\Gamma}$ and $\boldsymbol{\Theta}$ (resp. $\boldsymbol{\theta}$ and $\boldsymbol{\eta}$), making optimization challenging. To tackle this challenge, we use the bilinear relaxation formula introduced in Sec. 3 to relax these terms into linear ones. The first relaxation is:
@@
where in accordance with equation (3), the coefficients can be calculated as
@@
Here we use underline (resp. overline) to denote the lower (resp. upper) bounds of matrices $\boldsymbol{\Theta}$ and $\boldsymbol{\Gamma}$. It is also worth noting that the range $\underline{\boldsymbol{\Theta}}$, $\overline{\boldsymbol{\Theta}}]$ of $\boldsymbol{\Theta}$ can be derived from the range $\underline{\boldsymbol{\theta}},\overline{\boldsymbol{\theta}}]$ of $\boldsymbol{\theta}$ based on relation $\boldsymbol{\Theta}=\boldsymbol{\theta}\boldsymbol{\theta}^{\top}$, as explained in Sec. 5.3.
@@
where in accordance with equation (3), the coefficients can be calculated as
@@
By substituting back $\boldsymbol{\theta}\boldsymbol{\theta}^{\top}\Leftarrow\boldsymbol{\Theta}$, $\text{mat}(K\mathbf{B}_{2}\mathbf{p})\Leftarrow\boldsymbol{\Gamma}$ and $2\mathbf{A}\mathbf{p}\Leftarrow\boldsymbol{\eta}$, we get the final form of the lower bound function:
@@
1. Optimization over $\mathbf{p}$. This subproblem is a linear program in terms of the correspondence vector $\mathbf{p}$.
@@
2. Optimization over $\boldsymbol{\theta}$. This subproblem is a quadratic program in terms of the transformation vector $\boldsymbol{\theta}$.
@@ 5.2 Branching strategy
According to the result stated in Sec. 3, for the bilinear term $\boldsymbol{\theta}^{\top}\boldsymbol{\eta}$, for its linear lower bound function to converge to it, it suffices to branch over $\boldsymbol{\theta}$ and leave the range of $\boldsymbol{\eta}$ fixed. Here, utilizing $\boldsymbol{\eta}=-2\mathbf{A}\mathbf{p}$, the fixed range $\underline{\boldsymbol{\eta}},\overline{\boldsymbol{\eta}}]$ of $\boldsymbol{\eta}$ can be computed as
$\underline{\eta}_{i}=\min_{\mathbf{p}\in\Omega}(-2\mathbf{A}\mathbf{p})_{i},\quad\overline{\eta}_{i}=\max_{\mathbf{p}\in\Omega}(-2\mathbf{A}\mathbf{p})_{i}$ (16)
Likewise, for the bilinear term $\text{trace}(\boldsymbol{\Gamma}\boldsymbol{\Theta})$, for its linear lower bound function to converge to it, we only need to branch over $\boldsymbol{\theta}$ (since the range of $\boldsymbol{\theta}$ determines the range of $\boldsymbol{\Theta}$ as explained in Sec. 5.3) and leave the range of $\boldsymbol{\Gamma}$ fixed. Here, utilizing $\boldsymbol{\Gamma}=\text{mat}(K\mathbf{B}_{2}\mathbf{p})$, the fixed range $\underline{\boldsymbol{\Gamma}},\overline{\boldsymbol{\Gamma}}]$ of $\boldsymbol{\Gamma}$ can be computed as follows:
$\underline{\Gamma}_{ij}=\min_{\mathbf{p}\in\Omega}[\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})]_{ij},\quad\overline{\Gamma}_{ij}=\max_{\mathbf{p}\in\Omega}[\text{mat}(\mathbf{K}\mathbf{B}_{2}\mathbf{p})]_{ij}$ (17)
With these choice of $\underline{\boldsymbol{\Gamma}}$ and $\overline{\boldsymbol{\Gamma}}$, it has been empirically demonstrated that
@@
Please refer to Sec. 5.6 for relevant experiments and A for a weaker assertion. This result implies that problem (15a) (15b) is now a convex quadratic program. Consequently, utilizing solvers for convex quadratic programs, such as MATLAB’s quadprog function, guarantees global optimality. We will address the proof of (18) in our future research.
We also note that (16) and (17) are $n_{p}$-cardinlaity linear assignment problems, which can be efficiently solved by the aforementioned algorithms.
@@ 5.3 The range of $\Theta$ from the range of $\theta$
$\underline{\Theta_{ij}},\overline{\Theta_{ij}}]=\begin{cases}[\underline{\theta_{i}},\overline{\theta_{i}}]\cdot[\underline{\theta_{j}},\overline{\theta_{j}}],&\text{if}\ i\neq j\\ [\underline{\theta_{i}},\overline{\theta_{i}}]^{2},&\text{if}\ i=j\end{cases}$
@@
$\underline{\theta_{i}},\overline{\theta_{i}}]\cdot[\underline{\theta_{j}},\overline{\theta_{j}}]=[\min(\underline{\theta_{i}}\underline{\theta_{j}},\underline{\theta_{i}}\overline{\theta_{j}},\overline{\theta_{i}}\underline{\theta_{j}},\overline{\theta_{i}}\overline{\theta_{j}}),\max(\underline{\theta_{i}}\underline{\theta_{j}},\underline{\theta_{i}}\overline{\theta_{j}},\overline{\theta_{i}}\underline{\theta_{j}},\overline{\theta_{i}}\overline{\theta_{j}})]$
@@
[…]
@@ 5.5 Branch-and-Bound
With the preceding preparations in place, we now apply the Branch-and-Bound (BnB) algorithm to optimize the objective function $E$. We begin by forming an initial hypercube $M$ using the initial range of the parameters $\boldsymbol{\theta}$.
In each iteration, the algorithm selects the hypercube with the lowest lower bound for subdivision. This process progressively tightens the global lower bound. Simultaneously, the upper bound is refined by evaluating $E(\mathbf{p})$ using solutions derived during the lower bound computation. For a complete overview of the process, see Algorithm 1.
@@ 5.6 Positive Semidefiniteness of $\mathbf{H}^{0}+\mathbf{C}$
To verify the positive semidefiniteness of the matrix $\mathbf{H}^{0}+\mathbf{C}$ ($\mathbf{H}^{0}+\mathbf{C}\succeq 0$), we conducted an experiment using the separate outliers and inliers test, as detailed in Section 8.1.1. The outlier-to-data ratio was set at 0.3.
Figure 3 displays histograms of the smallest eigenvalues of $\mathbf{H}^{0}+\mathbf{C}$ across 100 different registration instances. The results consistently show that the smallest eigenvalue is non-negative, providing strong evidence that $\mathbf{H}^{0}+\mathbf{C}$ is indeed positive semidefinite. This outcome highlights the reliability and effectiveness of the proposed algorithm.
@@
5.7 BnB Algorithm Convergence
We assessed the convergence of our BnB algorithm using the separate outliers and inliers test (Sec. 8.1.1) with an outlier-to-data ratio of $0.3$. The results in Figure 4 reveal two key points:
  1. 1.
    A smaller initial range for $\boldsymbol{\theta}$ yields a tighter duality gap, which improves both upper and lower bounds. This suggests that leveraging a tighter initial range can speed up convergence.
  2. 2.
    Simpler problem instances (e.g., the fish test vs. the character test) have a tighter duality gap, leading to faster convergence.
Based on observation 2, a fixed duality gap threshold is an unsuitable stopping criterion for all problem types, as it can cause either excessive computation time or premature termination. Therefore, consistent with [5], we use the maximum branching depth as our stopping criterion.